OpenStudy (anonymous):

what is (3+2*6 to the ^2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@saxqueen352 Are you still there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(3 + 2*6)^2 = (3+12)^2 = 15^2 = 15*15 = 225

OpenStudy (anonymous):

can you try to be more specific

OpenStudy (anonymous):

no, this is the most specific you could be with this question? what do you not understand??

OpenStudy (skullpatrol):

@aerospacenerd please don't give away the answer.

OpenStudy (anonymous):

nothing its fine just got answer